OpenStudy (anonymous):

No clue on this one: "Find the sum, if it exists: 125-25+5-1+1/5..."

OpenStudy (anonymous):

there are no notations on when the sequence terminates, just the ... at the end. I can see the sequence is being divided by -5, but...?

OpenStudy (anonymous):

\[a= 125\] \[r =\frac{-1}{5}\]

OpenStudy (anonymous):

\[S = \frac{125}{1 + \frac{1}{5}}\]

OpenStudy (anonymous):

\[S = \frac{625}{6}\]

OpenStudy (anonymous):

i'm confused - the ... means the sequence doesn't terminate though, or?

OpenStudy (anonymous):

\[S = \frac{a(1 -r^n)}{(1-r)}\]

OpenStudy (anonymous):

when r<1 and n-> infinity r^n ~0 Hence the formula becomes \[S = \frac{a}{1-r}\]

OpenStudy (anonymous):

\[\S_{\infty}=a1/(1-r)\] is the formula i have, so 125/1.2 = 104.167?

OpenStudy (anonymous):

r=-1/5 so denominator becomes 1 +1/5 =6/5 = 1.2 yeah you are correct